The following is a hypothetical example. Wheat production in Namibia rose dramatically in 2018 due to favorable farming conditions. in 2017 quantity demanded was 2.66 million and average price (per bushel) was N$4.70 and in 2018 Quantity demanded was 2.82 million and average price (per bushels) was N$3.72 (a) Using the arc elasticity (midpoint) formula, calculate the price elasticity of demand for wheat.


1
Expert's answer
2021-04-22T07:53:48-0400

the arc elasticity of demand is given as

"E^{d}=-\frac{\Delta{Q}}{\Delta{P}}×\frac{P_{1}+P_{2}}{Q_{1}+Q_{2}}"

Where "\Delta{Q}" is given as $0.16 ($2.82-$2.66) and "\Delta{P}" is given as -$0.98 ($3.72-$4.70)

"P_1" is given as $4.70 and "P_2" as $3.72


"E^{d}=-\frac{0.16}{-0.98}×\frac{4.70+3.72}{2.66+2.82}"


"E^d=0.25" (demand is inelastic)

The price elasticity of demand for wheat is 0.25 which is inelastic